OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 6x - 3y = 12. slope = 2; y intercept at (0, -4) slope = -2; y intercept at (0, 4) slope = one half; y intercept at (0, 15) slope = one half; y intercept at (0, 9)

OpenStudy (anonymous):

A

OpenStudy (anonymous):

When in standard form: m (slope) = -A/B b (y-int) = C/B

OpenStudy (anonymous):

Standard Form: Ax + By = C

OpenStudy (anonymous):

-3y=-6x+12 y=2x+12 so it's A